OpenStudy (anonymous):

f(x)=(1-x-9)/(x-1)(x-3)=A/(x-1)+B/(x-3) find a and b

OpenStudy (anonymous):

a = 9/2 , b = -11/2

OpenStudy (anonymous):

Thank you agian BRO! I only have a few more then im finally done with this huge papper

OpenStudy (anonymous):

express the RHS as (A(x-3)+B(x-1)) / (x-1)(x-3) this means that Ax-3A +Bx-B = 1-x-9 = -x-8 this means that -3A-B = -8 and A+B = -1